### v1

# Training objective drives the consistency of representational similarity across datasets

Abstract
The Platonic Representation Hypothesis claims that recent foundation models are converging to a shared representation space as a function of their downstream task performance, irrespective of the objectives and data modalities used to train these models [35]. Representational similarity is generally measured for individual datasets and is not necessarily consistent across datasets. Thus, one may wonder whether this convergence of model representations is confounded by the datasets commonly used in machine learning. Here, we propose a systematic way to measure how representational similarity between models varies with the set of stimuli used to construct the representations. We find that the objective function is the most crucial factor in determining the consistency of representational similarities across datasets. Specifically, self-supervised vision models learn representations whose relative pairwise similarities generalize better from one dataset to another compared to those of image classification or image-text models. Moreover, the correspondence between representational similarities and the models’ task behavior is dataset-dependent, being most strongly pronounced for single-domain datasets. Our work provides a framework for systematically measuring similarities of model representations across datasets and linking those similarities to differences in task behavior.
1 Introduction
Representation learning has seen remarkable progress in recent years, with state-of-the-art models achieving or even surpassing human-level performance in a wide range of computer vision tasks [65, 18, 64, 85, 61]. The claim of The Platonic Representation Hypothesis is that with this progress in task performance, model representation spaces tend to approach each other irrespective of the models’ training data or architecture [35]. Even between models trained on different modalities, better downstream task performances appear to result in more similar representations. Two questions naturally arise from this phenomenon: Do pairwise similarities of model representations transfer from one dataset to another, and are they generally linked to the models’ downstream behavior?
Investigating the correspondence between representation and behavior has a long history in representation learning and adjacent fields [e.g. 32, 74]. Two models can have very different intermediate layer representations even though they show the same task behavior [32, 60, 48]. However, if two models show different behavior, one can be certain that the output layer representations are different, because identical output representations map to the same behavior and, vice versa, different output layer representations map to different behavior. What remains unclear is how behavior (e.g., downstream task accuracy) is affected by the convergence of model representations (i.e., the representations are similar according to a chosen similarity measure). Does this relationship depend on the specific similarity measure at hand, or may it be determined by the nature of a dataset?
The field lacks consensus on how to define (pairwise) representational similarity [cf. 45, 42, 81, 22, 74] and how to systematically measure its relationship to behavior [cf. 26, 60, 74, 48]. Here, we use similarity measures that are widely used in the community and have been empirically proven to be useful for measuring representational similarity of neural networks, such as CKA [42, 67] or RSA [45, 44]. However, pairwise representational similarity is usually measured for a single dataset and not across datasets. It may be the case that the representations of two vision models are highly similar for a dataset of raccoons but very different for a dataset of sunflowers — due to their different domains. Similarly, the representations of two models may exhibit a high similarity value for satellite data but can show a low similarity for images of fruit — because of the images’ different structure.
Thus, it appears crucial to scrutinize the factors that determine pairwise representational similarity across sets of stimuli. Knowing about those factors will enable us to generalize pairwise representational similarity across datasets and more effectively find model sets that have learned a similar representation of the world. For a large set of diverse vision models, we examine their pairwise representational similarities for both structured and unstructured datasets from different domains, using the publicly available Visual Task Adaptation Benchmark [VTAB; 87]. Our contributions and findings are as follows:
  • We propose a principled way of measuring the consistency of pairwise similarity across datasets, by measuring whether the (relative) pairwise similarities of model representations from one dataset are transferable to the (relative) pairwise similarities of another dataset.
  • We find that the objective function is the most crucial factor for determining the consistency of pairwise representational similarities across datasets, whereas architecture, model size, and training data do not appear to be qualitatively meaningful factors.
  • Pairwise similarities in SSL models generalize more reliably across stimulus sets compared to image-text and supervised models, which show high variance in their consistency due to dataset dependence.
  • We demonstrate that our findings hold irrespective of whether the kernel in the similarity measure emphasizes local or global representational structure.
  • The pairwise similarities of model representations show a strong correlation with the models’ differences in task performance for single-domain datasets, while multi-domain datasets show highly variable and specialized datasets consistently low correlations.
2 Related work
Measuring the similarity of deep neural networks. Representational similarity measures aim to quantify to what extent the representations of two models corresponding to the same input overlap. In the field of machine learning, they have been used as a tool to understand deep neural networks [51, 63, 56], investigate human-machine alignment [74, 82], and as an objective for model distillation [79, 70, 88]. Various ways of measuring representational similarity have been proposed, including Canonical Correlation Analysis [CCA; 34, 66, 58], Singular Vector Canonical Correlation Analysis [SVCCA; 66], Centered Kernel Alignment [CKA; 42, 14], and Representational Similarity Analysis [RSA; 45], among others [e.g. 81, 20, 5, 16]. They can be distinguished from functional similarity measures, such as performance difference [39], error consistency [25], and true-positive agreement [29], which are based on model outputs rather than internal representations [39], or model stitching [4, 15, 57], which measures compatibility, not similarity, of internal representations [33].
Importantly, similarity measures may differ in the invariances they evoke and the properties of representational spaces they capture [39]. For example, local structure may be better captured by measures based on nearest neighbors [35] or localized kernels [42], while other kernels may better capture global similarity. Some argue that similarity measures have to satisfy certain conditions, such as the triangle inequality and symmetry to qualify as measures of similarity [e.g. 81, 23, 30, 37], while others see value in relaxing the conditions of similarity [e.g. 9, 10, 74, 48]. Here, we use CKA, as it allows probing both local and global representation structures by varying the kernel, does not require the compared representations to be of the same dimension, and is widely used in the research community [e.g. 68, 55, 70, 88, 21, 6].
Drivers of similarity in vision models. In contrast to language processing systems [e.g. 8, 1, 76], computer vision models are trained with a wide range of different learning objectives. These include (a) supervised learning objectives using classification datasets such as ImageNet-1k [46], or ImageNet-21 [69]; (b) SSL objectives with a reconstruction or multi-view matching loss [12, 84, 11, 31]; and (c) objectives for jointly learning image and text representations [65, 36, 75, 86]. This makes them interesting candidate models for analyzing representational similarities.
The trend to larger model sizes [e.g. 18, 1] combined with increased dataset sizes [e.g. 72, 64] and multi-modality [e.g. 65, 75, 86] appears to be leading to a convergence of model representations [35]. Similarity between otherwise identical networks trained from different random initializations increases with width [59, 42], as predicted by theoretical studies of properties of neural networks in the infinite width limit [50, 17, 83]. Huh et al. [35] hypothesize that convergence is further driven by task generality and inductive bias. Other works find that architecture and dataset [68] play a central role in determining the representation structure of a (pre-)trained model.
Consistency of similarity. In addition to the similarity function itself, any comparison of representations depends on the layer from which the representations have been extracted [42, 68, 7] and the dataset for which the representations have been computed [16, 38, 7]. For example, measuring similarity for a subset of a data distribution can yield different similarities than measuring it for the whole dataset, albeit maintaining the ranking of similarities between models [7]. From a behavioral perspective, it has been shown that the ranking of the agreement of ImageNet classifiers pairs generalizes to out-of-distribution data [3]. Due to the limited scope of prior work, such as the inclusion of only a small number of models [38, 7], the driving factors of the consistency of representational similarities across vision datasets have not yet been identified.
3 Methods
Figure 1: Pairwise similarity analysis framework. Let ${\bm{A}}$ and ${\bm{B}}$ be two sets of stimuli and $\Theta$ and $\Phi$ be two sets of models. For each dataset, we extract representations from all functions within the two model sets. Subsequently, for each pair of models $(f_{\theta},f_{\phi})$, where $f_{\theta}\in\Theta$ and $f_{\phi}\in\Phi$, we compute the CKA similarities between their representations, yielding a similarity vector ${\bm{s}}_{\left({\bm{A}},\Theta,\Phi\right)}$. This vector can be displayed as a matrix, where each entry represents the similarity between two models for a single dataset. In the scatter plot, we contrast two such vectors computed on the same models but evaluated on different datasets. The Pearson correlation coefficient $\rho$ between the similarities quantifies the consistency of similarities across the two datasets. The distribution of $\rho$ across all dataset pairs indicates the stability of (relative) representational similarities across stimuli.
In this section, we propose a novel approach for analyzing the consistency of representational similarities across multiple datasets. Fig. 1 provides a schematic overview of this framework.
Extracting representations. We are interested in comparing the similarities of different vision models. We perform this comparison by extracting their latent representations and examining them. Let $f_{\theta}:\mathbb{R}^{d}\mapsto\mathbb{R}^{p}$ be a pretrained neural network function parameterized by a fixed set of parameters $\theta$ that maps the $d$-dimensional images to $p$-dimensional vector representations. Let ${\bm{X}}\in\mathbb{R}^{n\times d}$ be a dataset of $n$ stacked images. For each image ${\bm{x}}\in\mathbb{R}^{d}$ in a dataset, we extract its corresponding latent representations $f_{\theta}\left({\bm{x}}\right)={\bm{z}}\in\mathbb{R}^{p}$. For supervised models, we typically use the penultimate layer to extract this representation; for self-supervised vision models, the average pooling layer; and for image-text models, the image encoder. We denote by ${\bm{Z}}\in\mathbb{R}^{{n}\times p}$ the matrix of stacked vector representations, which is unique for each model and dataset combination.
Computing representational similarities for a dataset. We compute the similarity between two models $f_{\theta}$ and $f_{\phi}$ for a dataset by applying CKA [14, 42] to their vector representations ${\bm{Z}}^{f_{\theta}}$ and ${\bm{Z}}^{f_{\phi}}$. CKA computes the similarity based on the normalized Hilbert-Schmidt Independence Criterion [28], applied to the kernel matrices of both representations. Using CKA with a linear kernel focuses on global similarity structure, while an RBF kernel with small $\sigma$ measures local similarity structure [42, 2]. We compared their behavior (shown in Fig. 3 and in Appx. E) and observed similar trends for both kernels. Therefore, we use CKA with a linear kernel for the remainder of this paper if not mentioned otherwise.
Computing the similarity requires the full kernel matrix over the dataset, which scales quadratically with the number of data points. Hence, the exact computation becomes intractable for large datasets. However, we demonstrate in Appx. C that a diverse subset of $10{,}000$ samples is sufficient to accurately estimate the similarities of two models for an ImageNet-scale dataset.
Representational similarities across datasets. Because image representations depend on both a model and a dataset, the representational similarities between two models may vary across different datasets. To investigate this, we analyze the consistency of pairwise model similarities across multiple datasets. We compute the correlation coefficient for pairwise model similarities (measured for a large set of model pairs) between two datasets (i.e., can the pairwise model similarities for one dataset be expressed as a function of the similarities for a different dataset). In addition, we partition the models into specific sets determined by the models’ training factors.
Specifically, we scrutinize how the pairwise similarities of models from two model sets, $\Theta$ and $\Phi$, correspond between two datasets, ${\bm{A}}$ and ${\bm{B}}$. By computing the pairwise representational similarities between all model pairs $(f_{\theta},f_{\phi})$, separately for each dataset, where $f_{\theta}\in\Theta$, $f_{\phi}\in\Phi$ and $f_{\theta}\neq f_{\phi}$, we obtain a similarity vector ${\bm{s}}\in\mathbb{R}^{k}$, where $k$ denotes the number of distinct model combinations. A similarity vector for two sets of models and a dataset ${\bm{X}}$ is then defined as follows:
${\bm{s}}_{\left({\bm{X}},\Theta,\Phi\right)}=\left(\mathrm{CKA}({\bm{Z}}^{f_{\theta}}_{\bm{X}},{\bm{Z}}^{f_{\phi}}_{\bm{X}})\mid f_{\theta}\in\Theta,f_{\phi}\in\Phi,f_{\theta}\neq f_{\phi}\right)^{\top}\in\mathbb{R}^{k}.$
To quantify the consistency of similarities between two datasets ${\bm{A}}$ and ${\bm{B}}$, we use the Pearson correlation coefficient between the similarity vectors ${\bm{s}}_{({\bm{A}},\Theta,\Phi)}$ and ${\bm{s}}_{({\bm{B}},\Theta,\Phi)}$, i.e., $\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)$. The Pearson correlation measures the degree to which the similarity trends between models are preserved across datasets, focusing on the relative positioning of model pairs rather than the absolute similarity values. This correlation measures how well the models’ representational similarities are preserved across datasets. A high Pearson correlation indicates that the ordering of model similarities of two model pairs remains stable between datasets. In other words, if two models $f_{\phi}$ and $f_{\theta}$ are more similar than $f_{\phi}$ and $f_{\psi}$ on one dataset, a high Pearson correlation coefficient indicates that this relationship transfers to the other dataset. Conversely, a low correlation value suggests that the models’ relationships shift significantly depending on the dataset, showing variability in processing different data distributions. Thus, the Pearson correlation coefficient is a meaningful indicator of how dataset-dependent or dataset-invariant the model similarities are.
To assess the consistency of these correlations across dataset pairs, we examine the distribution of correlation coefficients for all possible dataset pairs within the set of available datasets $\mathcal{D}$:
${\bm{R}}(\Theta,\Phi)=\left\{\rho\left({\bm{s}}_{({\bm{A}},\Theta,\Phi)},{\bm{s}}_{({\bm{B}},\Theta,\Phi)}\right)\mid{\bm{A}},{\bm{B}}\in\mathcal{D},{\bm{A}}\neq{\bm{B}}\right\}.$
The distribution of the correlation coefficients across all dataset pairs indicates the consistency of the pairwise model similarities.
4 Experiments
Our experiments investigate the consistency of representational similarities between vision models across various datasets using our proposed framework. Following the experimental setup (Sec. 4.1), we first analyze how model similarities transfer across datasets (Sec. 4.2 and 4.4) and find consistent model groups through similarity clustering (Sec. 4.3). We then investigate how model and dataset characteristics influence similarity consistency (Sec. 4.5 and 4.6), and conclude by examining the relationship between representational similarities and models’ performance differences on downstream tasks (Sec. 4.7). All code to run our analyses and reproduce the results presented in this section is publicly available on GitHub: https://github.com/lciernik/similarity_consistency.
4.1 Setup
Models. We evaluated 64 vision models, representing a diverse range of architectures, training paradigms, and parameter scales. A complete list of models and their characteristics can be found in Appx. A. We used the Python package thingsvision [62] to extract the model representations.
Grouping models by attributes. Instead of focusing on the representational similarities of individual model pairs, we analyze the aggregated similarities between sets of models. We categorize models into sets based on four attributes: training objective, training data, model architecture, and model size (see Tab. 1). Grouping models by these attributes allows us to examine how different model characteristics relate to the behavior of representational similarities while acknowledging that these attributes are often correlated across models. For example, we measured the similarity between models trained with different objectives (e.g., supervised, SSL, and image-text) and assessed the consistency of representational similarities within and between each model group.
Table 1: Number of evaluated models in each attribute category.
Category Nr. models Category Nr. models
Training objective Training data
Image-Text (Img-Txt) 14 ImageNet-1k (IN1k) 37
Self-Supervised (SSL) 20 ImageNet-21k (IN21k) 9
Supervised (Sup) 30 Large 11
XLarge 7
Architecture Class Model Size
Convolutional (CNN) 24 small $<100M$ parameter 32
Transformer (TX) 40 medium $<200M$ parameter 14
large $<400M$ parameter 8
xlarge $>400M$ parameter 10
Datasets. We evaluated pairwise model similarities across 20 classification datasets from the CLIP benchmark  [^1] https://github.com/LAION-AI/CLIP_benchmark and 3 datasets from Breeds [71]. This set includes various VTAB datasets as well as ImageNet-1k. The CLIP benchmark converts multi-label to single-label classification datasets by cropping images to the bounding box of each object. Following the categorization proposed in [87], we classified the datasets into three main types: natural (e.g., ImageNet-1k), specialized (e.g., PCAM), and structured (e.g., DTD) image datasets. Furthermore, we partition the natural image datasets into single-domain and multi-domain categories. A list of all datasets and their categorization can be found in Tab. 3.
4.2 Do representational similarities transfer across datasets?
The transferability of model similarities across diverse datasets is crucial for model comparison and selection. If these similarities are consistent across different datasets, it will enable us to evaluate models on a single dataset and generalize their relationships to other datasets. We analyzed this by computing pairwise CKA similarities for all model pairs across our dataset collection.
Model similarities (e.g., the range of values and the similarity trends) appear to vary across datasets (Fig. 2). We observe the highest similarity within image-text models (yellow boxes), while a group of self-supervised models (white boxes) shows consistent similarity trends across datasets. Yet, there are also model groups with higher variability (cyan boxes), mainly containing supervised models. The standard deviation matrix (rightmost matrix in Fig. 2) quantifies these differences, indicating that model similarities do not transfer uniformly across datasets.
To further investigate the transferability of model similarities across datasets, we examine the relationship between the mean and standard deviation of similarity values. The left panel in Fig. 3 shows an inverse U-shape trend (further analyzed in Appx. D): model pairs with extreme mean similarities exhibit low variability, while variability for mid-range similarities is higher. To verify that this observed variability is not an artifact of the similarity metric, we compare the mean similarity values obtained using linear CKA with two other metrics: CKA with an RBF kernel where $\sigma=0.4$ (a local similarity measure) and RSA using Spearman correlation (a global similarity measure, Fig. 3). The high correlations among these measures confirm that mid-range similarity values consistently vary across metrics, indicating limited stability across datasets.
Given this variability, we investigate persistent trends in model relationships across datasets by addressing two questions: a) Can we identify subgroups of models whose similarity to other models (consistently) cluster across different datasets? b) Do the relative representational similarities within or between these subgroups remain stable/consistent across datasets? Answering these questions affirmatively would allow us to make generalizable, albeit potentially weaker, claims about models’ relationships from a single data source.
[Uncaptioned image] Figure 2: Matrices of pairwise representational similarities using linear CKA. Left to right: natural multi-domain, natural single-domain, and specialized datasets, followed by average and standard deviation matrices across all datasets. Models are ordered by a hierarchical clustering of the mean similarity matrix. Boxes highlight more stable (yellow, white) and variable (cyan) groups of model pairs across the datasets.
[Uncaptioned image] Figure 3: Variability of representational similarities across datasets. Left: Mean versus standard deviation of pairwise similarities. Center/Right: Mean similarity values using CKA RBF ($\sigma=0.4$) or RSA as a function of linear CKA. High Pearson correlation coefficients indicate consistency across different similarity metrics.
4.3 Do representational similarities cluster according to model categories?
We qualitatively assess the alignment between our predefined model categories (Sec. 4.1) and the clustering patterns observed in the t-SNE embedding of similarity matrices, as illustrated in Fig. 4. The training objective appears to be the most distinctive categorization, with all image-text and most SSL models forming distinct clusters across all three datasets. The training objective and data categories show substantial overlap, potentially confounding further analysis. For instance, only image-text models have been trained on XLarge datasets, while all models trained on ImageNet-21k are supervised models (Appx. A). Additionally, no image-text model has been trained on ImageNet-1k or ImageNet-21k. In contrast, no obvious clustering is discernible based on architecture or model size. These observations provide initial evidence that the training objective might play a key role in representational similarity.
Figure 4: t-SNE visualization of model similarity matrices for ImageNet-1k, Flowers, and PatchCamelyon (PCAM) datasets. Embeddings are color-coded by model attributes: Training objective, Architecture, Training data, and Model size.
4.4 Do relative representational similarities remain consistent across different datasets?
Given the observation that representational similarities are not directly transferable across datasets but depend on both the models and the dataset, we now examine the consistency of relative representational similarities across dataset pairs. Following the method described in Sec. 3, the top row of Fig. 5 shows the similarities of all model pairs across three dataset pairs, showing positive correlations. Across all dataset pairs, the mean Pearson correlation coefficient is 0.756 (std=0.124). This indicates that the relative ordering of pairwise model similarities is largely consistent across datasets, suggesting a degree of transferability in the relative relationships of representational similarities. However, the high standard deviation and the distributions of similarities shown in Fig. 5 suggest the existence of model-pair groups that may exhibit distinct consistencies of similarity values.
The second row in Fig. 5 shows model pairs within the same training objective, highlighting differences in group-specific consistency levels. We find that SSL models show the strongest similarity consistency across datasets, image-text models form a cluster of generally high similarities, and supervised models achieve the weakest consistency. Next, we will analyze the aggregated measure across all dataset pairs and revisit all categories in the following section.
[Uncaptioned image] Figure 5: Pairwise similarities across ImageNet-1k, Flowers, and PCAM datasets. Both rows show representational similarities for three dataset pairs (ImageNet-1k vs. Flowers, ImageNet-1k vs. PCAM, Flowers vs. PCAM). Each point represents a model pair, with its position indicating similarity scores in the two compared datasets. Top row: Comparison with Pearson correlation coefficients for all model pairs. Bottom row: Same data, but colored points highlight model pairs within the same training objective category (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$, ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, and ${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$). Gray points represent the remaining cross-category pairs.
4.5 Which model categories influence relative similarity consistency?
Fig. 6 shows the distribution of Pearson correlation coefficients ${\bm{R}}(\Theta,\Phi)$ for each combination of model sets $(\Theta,\Phi)$ across all dataset pairs.
Training objective. Self-supervised models (${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$) show the highest relative similarity consistency, even when compared to models with different training objectives. This suggests that their relative similarities are most transferable across datasets, likely due to the ability of SSL to capture dataset-independent features. In contrast, supervised models show the weakest correlations, particularly when compared to image-text models (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$), indicating a less reliable transfer of relative pairwise similarities across datasets.
Model architecture. While no significant differences in similarity consistency exist between transformers and convolutional networks, comparisons across architectures show slightly lower consistency, potentially indicating distinct architectural inductive biases.
Training data. We find a low median correlation and a high variance for the model set pairs $({\color[rgb]{1,0.37,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.37,1}\pgfsys@color@cmyk@stroke{0}{0.63}{0}{0}\pgfsys@color@cmyk@fill{0}{0.63}{0}{0}\Phi_{\text{IN21k}}},{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@cmyk@stroke{0}{0}{0}{0.50}\pgfsys@color@cmyk@fill{0}{0}{0}{0.50}\Phi_{\text{Large}}})$ and $({\color[rgb]{1,0.37,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.37,1}\pgfsys@color@cmyk@stroke{0}{0.63}{0}{0}\pgfsys@color@cmyk@fill{0}{0.63}{0}{0}\Phi_{\text{IN21k}}},{\color[rgb]{0.74,1,0.24}\definecolor[named]{pgfstrokecolor}{rgb}{0.74,1,0.24}\pgfsys@color@cmyk@stroke{0.26}{0}{0.76}{0}\pgfsys@color@cmyk@fill{0.26}{0}{0.76}{0}\Phi_{\text{XLarge}}})$. However, these effects may be confounded by the training objectives, as both ${\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}\pgfsys@color@cmyk@stroke{0}{0}{0}{0.50}\pgfsys@color@cmyk@fill{0}{0}{0}{0.50}\Phi_{\text{Large}}}$ and ${\color[rgb]{0.74,1,0.24}\definecolor[named]{pgfstrokecolor}{rgb}{0.74,1,0.24}\pgfsys@color@cmyk@stroke{0.26}{0}{0.76}{0}\pgfsys@color@cmyk@fill{0.26}{0}{0.76}{0}\Phi_{\text{XLarge}}}$ contain mainly image-text models, while ${\color[rgb]{1,0.37,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.37,1}\pgfsys@color@cmyk@stroke{0}{0.63}{0}{0}\pgfsys@color@cmyk@fill{0}{0.63}{0}{0}\Phi_{\text{IN21k}}}$ consists solely of supervised models. On the other hand, ${\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}}$, which includes models trained on a dataset similar to ImageNet-21k but with diverse training objectives, does not exhibit the same low consistency. Consequently, training data appears to have a less significant impact on the consistency of similarities than other factors.
Model size. Models within the same size category demonstrate higher consistency of similarity than models of different size categories (e.g., ${\color[rgb]{0.12,0.98,0.64}\definecolor[named]{pgfstrokecolor}{rgb}{0.12,0.98,0.64}\pgfsys@color@cmyk@stroke{0.86}{0}{0.34}{0.02}\pgfsys@color@cmyk@fill{0.86}{0}{0.34}{0.02}\Phi_{\text{small}}}$ and ${\color[rgb]{0.68,0.36,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0.36,1}\pgfsys@color@cmyk@stroke{0.32}{0.64}{0}{0}\pgfsys@color@cmyk@fill{0.32}{0.64}{0}{0}\Phi_{\text{xlarge}}}$). This indicates model size influences the transferability of relative similarities across datasets, though less significantly than the training objective.
Our analysis identifies the training objective as the primary factor influencing similarity consistency, with model size playing a less pronounced role. Self-supervised models consistently demonstrate above-average correlations across datasets. The effects of model architecture and training data are minimal, suggesting that these factors are less critical in determining the transferability of relative similarities across datasets (see Appx. F for an analysis using single-element model sets controlling for model architecture and size).
Figure 6: Distribution of correlation coefficients for model similarities across dataset pairs. Each subplot represents a model category: Training objective, Architecture, Training data, and Model size. Within each category, distributions show correlation coefficients between model set combinations (e.g., ${\bm{R}}($${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$)))). The boxes in each subplot are sorted in decreasing median correlation. The dotted line indicates the median, while the gray area spans the 25th to 75th percentiles of the correlation distribution without considering model categories.
4.6 Do dataset categories influence relative similarity consistency?
Previously, we analyzed similarity consistency from the perspective of models, noting substantial variation in consistency across all dataset pairs, where some pairs produce consistent representations while others do not. In this section, we shift our focus to identifying groups of datasets that exhibit more stable relative similarities, aiming to uncover variables that might explain these variations.
The left panel in Fig. 7 shows the Pearson correlations of all model pairs for each dataset pair individually, i.e., $\rho\left({\bm{s}}_{({\bm{A}},\Phi_{\text{All}},\Phi_{\text{All}})},{\bm{s}}_{({\bm{B}},\Phi_{\text{All}},\Phi_{\text{All}})}\right)$ for all datasets pairs $({\bm{A}},{\bm{B}})$. The consistency of representational similarities varies strongly across dataset pairs, showing higher consistency within dataset categories (especially natural multi-domain and structured datasets) than across them, as exemplified by high consistencies within natural multi-domain datasets but lower ones when paired with specialized datasets. Moreover, some datasets are inconsistent with almost any other dataset, particularly FGVC Aircraft and medical imaging datasets (Diabetic Retinopathy and PCAM). Interestingly, ImageNet-1k exhibits a milder yet significant pattern of inconsistency. Within the multi-domain category, this is especially pronounced for CIFAR-10 and CIFAR-100.
To investigate if specific dataset categories are responsible for the difference in consistency observed in the last section, we focus on the Pearson correlations for the model sets with the highest (${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$) and lowest (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) consistency (see the center and right panels of Fig. 7). The two matrices reveal distinctly different patterns, highlighting how model categories influence similarity consistency.
For SSL model pairs, we generally observe high consistency across most datasets, with FGVC Aircraft being an exception, showing lower correlations with the rest. This confirms that SSL models tend to produce more transferable similarity structures across datasets. In contrast, image-text/supervised model pairs show relatively higher consistency for multi-domain than single-domain datasets, suggesting better similarity transfer when semantic categories in the dataset match the training data of the models. However, Pascal VOC 2007 appears as an outlier when compared to ImageNet-1k and its subsets, likely because its bounding-box-cropped images lack contextual information, causing single-domain-like behavior.
Overall, we observe the highest consistency within multi-domain image datasets, suggesting that models are consistent when dealing with diverse semantic categories. In contrast, we find weaker consistency for (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) between multi-domain and single-domain datasets, indicating that even though these models effectively generalize across diverse datasets, their representational structures differ substantially when evaluated on single domains. Self-supervised models do not exhibit this sensitivity to datasets; their similarity relationships appear to be more stimuli-invariant, maintaining consistent similarities regardless of whether the data is multi-domain or single-domain.
[Uncaptioned image] Figure 7: Pearson correlation of similarities for all dataset pairs computed on different model sets. Left: all available model pairs; Center: SSL models (${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$); Right: image-text and supervised models (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$). Datasets are sorted by category and delineated by dotted lines.
4.7 Can representational similarity predict performance gaps?
[Uncaptioned image] Figure 8: Model similarity (CKA linear) vs. absolute difference in downstream task performance (top-1 accuracy) for model pairs across four dataset categories: natural multi-domain, natural single-domain, specialized, and structured. The color of each point indicates the downstream task accuracy of the better-performing model. Pearson correlation coefficients are shown in each subplot.
To investigate the relationship between downstream task performance and representational similarity, we correlate two measures for each dataset: (1) the CKA value for each model pair and (2) the absolute difference in classification performance, referred to as the performance gap (see Appx. B for details on performance computation). Fig. 8 shows this relationship across three datasets from each category. While prior work suggests well-performing models produce similar representations [35], we find this relationship is dataset-dependent.
Natural multi-domain datasets, such as ImageNet-1k, show weak negative correlation between similarity and performance gap. Many model pairs achieve high performance despite low representational similarity (yellow dots in the lower left corner). This suggests that multi-domain datasets might provide richer contextual information, enabling multiple high-performance strategies that do not converge to similar representations. However, correlation coefficients in this dataset category show notable variability (see Appx. G for further discussion).
Natural single-domain and Structured datasets, in contrast, exhibit a strong negative correlation between similarity and performance gap (see Fig. 15 for all datasets). The absence of high-performing, dissimilar model pairs suggests a close link between performance and representational similarity, indicating that successful models for these datasets rely on capturing specific discriminative features for closely related classes. Observing this behavior in both structured and single-domain datasets suggests that models may depend on representing more structural attributes to distinguish between classes. However, the limited number of structured datasets restricts broader conclusions.
In summary, our analysis shows varying trends depending on the dataset. Multi-domain datasets support a wider range of successful strategies, whereas performing well on single-domain and structured datasets appears to require a more limited set of features, resulting in stronger correlations between performance and similarity. This phenomenon aligns with findings from previous studies, which have shown that models can leverage contextual cues in multi-domain datasets to achieve strong performance without developing robust, generalizing features [49, 27, 24]. For specialized datasets, we find no or weak correlation between pairwise similarity and performance gap, along with reduced variation in both measures.
Our analysis provides clear evidence of variations in the relationship between performance gap and similarity across different datasets, highlighting the need for further research to identify the key factors driving this variation.
5 Discussion
The representations of the latest foundation models appear to converge to a canonical representation, irrespective of their training data and objectives [35]. However, this isotropy of model representations may just be an artifact of the datasets that the community commonly evaluates. Thus, in this paper, our goal was to present a systematic way of analyzing pairwise similarities of model representations across sets of stimuli. This approach allowed us to examine whether representational similarities generalize across datasets and identify variables influencing their consistency.
We found that the training objective of models is a central factor for determining the consistency of pairwise representational similarities across datasets, whereas architecture, model size, and training data appear less important. Among training paradigms, image-text training leads to the most similar representation spaces, irrespective of the models’ training data or architecture. For individual datasets, SSL model representations are not as similar among themselves as the representations of image-text models, but the pairwise similarities between SSL models and other models are highly consistent across different datasets.
We hypothesize that both image-text and supervised trained models learn distinct semantic categories during training that lead to representations whose features overfit to these categories and, therefore, do not generalize well to datasets that do not contain them. In contrast, self-supervised trained pure vision models—which are not constrained to learn explicit semantic categories during training [cf. 12, 40, 84, 11]—may develop more flexible representations that better accommodate diverse dataset stimuli. This leads to representations whose relative pairwise similarities vary less across datasets than the similarities of representations from image-text or supervised models. This finding supports our hypothesis that models trained on specific semantic categories develop dataset-dependent representations due to their features’ sensitivity to mismatches between training and evaluation categories.
An alternative explanation may be that the extracted representations from SSL models are further away from the models’ task behavior than the image encoder and penultimate layer representations from image-text and supervised models, respectively. Usually, SSL model representations are extracted from the average pooling layer [cf. 12, 13, 60, 78] which can be seen as further away from behavior than the penultimate layer of supervised models. One benefit of our framework is that it can be applied to the representations of any model layer, allowing future work to test this hypothesis. We note that, in contrast to image-text and supervised models, the tasks used to train SSL models are extremely heterogeneous. This may yet be another reason for the higher consistency of representational similarities among SSL models.
Finally, we have shown that the pairwise similarities between model representations are only in part predictive of their differences in downstream task behavior. While we observe a strong correlation on datasets with limited contextual information, for others, e.g., ImageNet-1k, there is no obvious correspondence between the pairwise similarities of model representations and the models’ differences in task performance. Moreover, the relationship between pairwise similarities and task performance is substantially different between natural multi-domain datasets, e.g., for ImageNet-1k vs. CIFAR-100 (see Fig. 8). This is surprising in light of recent findings that have shown a convergence of model representations as a function of task performance [35] and a strong linear relationship between a model’s ImageNet performance (i.e., performance on the training set distribution) and its downstream transfer accuracy [43]. This suggests that the transferability of pairwise similarities across data sets follows a different and more complex relationship than the transferability of task performance. While different objectives and hyperparameters lead to similar task performances [41], they seem to change the similarity spaces of the learned representations in a way that affects the relationship between pairwise model similarities and their differences in downstream behavior across datasets. Therefore, whether the claims of The Platonic Representation Hypothesis hold or not depends on the specific nature of a dataset rather than being generally true for all sets of stimuli used to compute model representations.
Conclusion.
In summary, we presented a blueprint for systematically analyzing pairwise representational similarities across different sets of stimuli. This framework allowed us to demonstrate that pairwise similarities rarely transfer from one dataset to another and depend on both the models’ objective functions and the datasets’ structure and domain. The community has long optimized for finding the optimal hyperparameters to improve task performance [e.g. 80, 73, 52, 19] but neglected its ramifications for the similarity structure of the models’ representation spaces. Only very few recent works investigated how the learned model features are affected by training tasks [e.g. 48, 41]. We hope that our work motivates to pinpoint the variables during model training that affect (dis-)similarities between model representations beyond the training dataset because identifying those variables will ultimately help to build more interpretable systems with which humans are more likely to interact [60, 47, 74, 61, 77].
Author Contributions
LC and LM managed the project, with LM providing overall research supervision. The codebase development was led by LC, with major contributions from LL and MM. Data gathering and preprocessing were performed by LC, LL, and JD. LC, LL, and MM created the figures and tables, incorporating valuable feedback from the other authors. The original draft of the manuscript was written by LC, LL, MM, JD, and LM, with SK providing critical revisions. All authors participated in the analysis and discussion of the results and contributed to the final manuscript.
Acknowledgments and Disclosure of Funding
LC is funded by the Hector Fellow Academy and is grateful for their support.
Appendix / Supplementary Material
Appendix A Models and Datasets
We considered 64 pretrained vision models (Tab. 2) and 23 downstream datasets (Tab. 3) in our analysis. For each dataset, we selected the training split, potentially subsetted for large datasets as described in Appx. C, to compute the model representational similarities, while we use the validation split to compute downstream performance.
Model name Rep. Layer Training objective Training data Training data class Architecture Architecture class Model size Model size class
Kakaobrain-Align pooler Img-Txt COYO-700M Large EfficientNet CNN 62.1M small
OpenCLIP-EVA01-g-14-plus-merged2b-s11b-b114k visual Img-Txt Merged2B XLarge ViT TX 1.4B xlarge
OpenCLIP-EVA01-g-14-laion400m-s11b-b41k visual Img-Txt LAION400M Large ViT TX 1.1B xlarge
OpenCLIP-EVA02-B-16-merged2b-s8b-b131k visual Img-Txt Merged2B XLarge ViT TX 149.7M medium
OpenCLIP-EVA02-L-14-merged2b-s4b-b131k visual Img-Txt Merged2B XLarge ViT TX 427.8M xlarge
OpenCLIP-RN50-openai visual Img-Txt WIT-400M Large ResNet CNN 102.0M medium
OpenCLIP-ViT-B-16-SigLIP-webli visual Img-Txt WebLI XLarge ViT TX 203.2M medium
OpenCLIP-ViT-B-16-laion2b-s34b-b88k visual Img-Txt LAION2B XLarge ViT TX 149.6M medium
OpenCLIP-ViT-B-16-laion400m-e32 visual Img-Txt LAION400M Large ViT TX 149.6M medium
OpenCLIP-ViT-B-16-openai visual Img-Txt WIT-400M Large ViT TX 149.6M medium
OpenCLIP-ViT-L-14-laion2b-s32b-b82k visual Img-Txt LAION2B XLarge ViT TX 427.6M xlarge
OpenCLIP-ViT-L-14-laion400m-e32 visual Img-Txt LAION400M Large ViT TX 427.6M xlarge
OpenCLIP-ViT-L-14-openai visual Img-Txt WIT-400M Large ViT TX 427.6M xlarge
vit-huge-patch14-clip-224.laion2b norm Img-Txt LAION2B XLarge ViT TX 632.1M xlarge
vit-large-patch14-clip-224.laion2b norm Img-Txt LAION2B XLarge ViT TX 304.0M large
dino-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
dino-vit-base-p16 norm SSL IN1k IN1k ViT TX 85.8M small
dino-vit-small-p16 norm SSL IN1k IN1k ViT TX 21.7M small
dino-xcit-medium-24-p16 norm SSL IN1k IN1k ViT TX 83.9M small
dino-xcit-small-12-p16 norm SSL IN1k IN1k ViT TX 25.9M small
dinov2-vit-base-p14 norm SSL LVD-142M Large ViT TX 86.6M small
dinov2-vit-giant-p14 norm SSL LVD-142M Large ViT TX 1.1B xlarge
dinov2-vit-large-p14 norm SSL LVD-142M Large ViT TX 304.4M large
dinov2-vit-small-p14 norm SSL LVD-142M Large ViT TX 22.1M small
jigsaw-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
mae-vit-base-p16 norm SSL IN1k IN1k ViT TX 86.4M small
mae-vit-huge-p14 norm SSL IN1k IN1k ViT TX 630.8M xlarge
mae-vit-large-p16 norm SSL IN1k IN1k ViT TX 303.3M large
mocov2-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
pirl-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
rotnet-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
simclr-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
swav-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
vicreg-rn50 avgpool SSL IN1k IN1k ResNet CNN 23.5M small
beit-base-patch16-224 norm Sup IN21k + IN1k IN21k ViT TX 86.5M small
beit-base-patch16-224.in22k-ft-in22k norm Sup IN21k IN21k ViT TX 102.6M medium
beit-large-patch16-224 norm Sup IN21k + IN1k IN21k ViT TX 304.4M large
beit-large-patch16-224.in22k-ft-in22k norm Sup IN21k IN21k ViT TX 325.8M large
convnext-base head.flatten Sup IN1k IN1k ConvNeXt CNN 88.6M small
convnext-large head.flatten Sup IN1k IN1k ConvNeXt CNN 197.8M medium
deit3-base-patch16-224 norm Sup IN1k IN1k ViT TX 86.6M small
deit3-base-patch16-224.fb-in22k-ft-in1k norm Sup IN21k + IN1k IN21k ViT TX 86.6M small
deit3-large-patch16-224 norm Sup IN1k IN1k ViT TX 304.4M large
deit3-large-patch16-224.fb-in22k-ft-in1k norm Sup IN21k + IN1k IN21k ViT TX 304.4M large
efficientnet-b3 avgpool Sup IN1k IN1k EfficientNet CNN 12.2M small
efficientnet-b4 avgpool Sup IN1k IN1k EfficientNet CNN 19.3M small
efficientnet-b5 avgpool Sup IN1k IN1k EfficientNet CNN 30.4M small
efficientnet-b6 avgpool Sup IN1k IN1k EfficientNet CNN 43.0M small
efficientnet-b7 avgpool Sup IN1k IN1k EfficientNet CNN 66.3M small
resnet152 avgpool Sup IN1k IN1k ResNet CNN 60.2M small
resnet50 avgpool Sup IN1k IN1k ResNet CNN 25.6M small
resnext50-32x4d global_pool Sup IN1k IN1k ResNeXt CNN 25.0M small
seresnet50 global_pool Sup IN1k IN1k SE-ResNet CNN 28.1M small
swin-base-patch4-window7-224 global_pool Sup IN1k IN1k Swin-Transformer TX 87.8M small
swin-base-patch4-window7-224.ms-in22k global_pool Sup IN1k IN1k Swin-Transformer TX 109.1M medium
swin-large-patch4-window7-224 global_pool Sup IN1k IN1k Swin-Transformer TX 196.5M medium
swin-large-patch4-window7-224.ms-in22k global_pool Sup IN1k IN1k Swin-Transformer TX 228.6M medium
vgg16 classifier.3 Sup IN1k IN1k VGG CNN 138.4M medium
vgg19 classifier.3 Sup IN1k IN1k VGG CNN 143.7M medium
vit-base-patch16-224 norm Sup IN1k IN1k ViT TX 86.6M small
vit-base-patch16-224.augreg-in21k norm Sup IN21k IN21k ViT TX 102.6M medium
vit-huge-patch14-224.orig-in21k norm Sup IN21k IN21k ViT TX 630.8M xlarge
vit-large-patch16-224 norm Sup IN1k IN1k ViT TX 304.3M large
vit-large-patch16-224.augreg-in21k norm Sup IN21k IN21k ViT TX 325.7M large
Table 2: Pretrained neural networks that we considered in our analyses.
Category Dataset Category Dataset
Natural Caltech101 Natural Cars
(multi-domain) CIFAR-10 (single-domain) FGVC Aircraft
CIFAR-100 Flowers
Country211 GTSRB
ImageNet-1k Pets
STL-10 SVHN
Pascal VOC 2007
Entity-30 (Breeds) Specialized Diabetic Retinopathy
Living-17 (Breeds) EuroSAT
NonLiving-26 (Breeds) PatchCamelyon (PCam)
RESISC45
Structured FER2013
Dmlab
DTD
Table 3: Datasets of clip-benchmark and breeds and their domains according to [87] with additional separation of natural images into multi- and single-domain.
Models. The models were chosen to represent a diverse set of training datasets, training objectives, model sizes, and architectures. Training objectives include standard supervised learning (Sup), self-supervised learning (SSL), and image-text alignment (Img-Txt). The supervised and SSL models were all pretrained on ImageNet-1k or ImageNet-22k, with the exception of DINOv2 models that were trained on the LVD-142M dataset. Image-text models were pretrained on image-text pairs ranging from 400M samples (e.g., LAION400M, WIT-400M) considered as Large to 2B samples (e.g., LAION2B) labeled as XLarge. In terms of architecture, the models employ a range of designs, including Convolutional Neural Networks (CNNs) such as ResNet, EfficientNet, ConvNeXt, and VGG, as well as Transformer-based models like ViT and Swin-Transformer. Further, we categorized the models into four different size buckets ranging from small, with around 12 million parameters, to xlarge, with over 1.4 billion parameters.
Preprocessing. To compute latent representations for each model/dataset combination, we used the Python package thingsvision [62]. Images were resized to 256px and center-cropped to 224px before applying the model-specific normalizations from the pretraining. Tab. 2 specifies exactly which layers were used for each model. For ViT models, we used the CLS token representations after the final LayerNorm module.
Appendix B Downstream task performance
We evaluated each model’s downstream task performance on every dataset by calculating the top-1 accuracy of a linear probe trained on features extracted from the corresponding training set. Each model was evaluated with 3 different random seeds on each dataset, and the mean top-1 accuracy across seeds was used. Hyperparameter selection was performed with a validation set of $20\%$ of the training set and the remaining training data for optimization. We followed the binary search procedure described in [65] and searched for the optimal weight decay parameter $\lambda$ in the interval between $10^{-6}$ and $10^{2}$ in 96 logarithmically spaced steps. This was done for all learning rates $\eta\in\{10^{i}\}_{i=1}^{4}$. After hyperparameter selection, the linear probe was retrained on the full training set and evaluated on the respective test set. All linear probes are trained for 20 epochs, using the AdamW optimizer [54] and a cosine schedule for learning rate decay [53].
Appendix C CKA sensitivity to number of samples in dataset
Our analysis of model similarities requires reliable CKA measurements while managing computational constraints. To identify the minimum number of samples needed for stable CKA values, we analyzed ImageNet-1k subsets with varying samples per class ($k\in{1,5,10,20,30,40}$). We generated pairwise similarity matrices (Fig. 9) and assessed stability by comparing the absolute differences between consecutive sample sizes (Fig. 10).
The results show that most CKA variants converged with 10,000 samples (k=10), with only CKA RBF ($\sigma=0.2$) requiring more samples for stability. Based on these findings, we set k=10 as our standard sample size and applied stratified subsampling to limit datasets larger than 10,000 samples. For all analyses, we selected CKA RBF ($\sigma=0.4$) and CKA linear as our local and global measures, respectively. While this evaluation used a slightly different set of models than our main analysis, it provides guidance about the number of samples that ought to be used to guarantee the stability of CKA measurements.
Figure 9: Pairwise model representational similarity matrices using different dataset sizes (1-40 samples per ImageNet-1k class) and metrics (CKA RBF 0.2-0.8 to linear). Rows indicate increasing samples per class, while columns progress from local to global similarity measures
Figure 10: Distribution of absolute differences of pairwise model representational similarities for each CKA variant measured on datasets with consecutive sample sizes per class ($1{\rightarrow}5$, $\cdots$, $30{\rightarrow}40$). Most CKA variants converge at $k=10$ samples per class, with CKA RBF ($\sigma=0.2$) requiring larger sample sizes for stability.
Appendix D Variation of CKA values near the upper bound
The CKA similarity measure is bounded between 0 and 1, with values close to 1 indicating highly similar representations. However, this upper bound might lead to saturation effects. To better differentiate between highly similar model representations, we apply two nonlinear transformations: $\arccos$ and $\tan$. Both transformations expand the range of high CKA values, with $\tan$ slightly reducing variation for low similarity values. These transformations amplify the inverse U-shape relationship noted in Section 4.2 (Figure 11), suggesting that the relationship between mean and variance of similarities is robust across different scales.
Figure 11: Mean versus standard deviation of CKA values after nonlinear transformations. Left: $\arccos$ transformation inverts the scale, mapping high similarities to low values and vice versa. Right: $\tan$ transformation. Both transformations make the inverse U-shape relationship more pronounced.
Appendix E Does local or global similarity differ?
As shown in the main text (Fig. 3), mean representational similarities across datasets demonstrate strong positive correlations when measured with different similarity metrics. In this section, we verify that our findings about similarity consistency remain valid when using both CKA RBF 0.4 and RSA Spearman metrics rather than just CKA linear.
To validate this stability across metrics, we correlate the consistency values across all dataset pairs and model set pairs. As shown in Fig. 12, there is a strong positive correlation between global (CKA linear) and local (CKA RBF with $\sigma=0.4$) measures, indicating that similarity consistency observations based on CKA linear can be generalized to the local measure. While the correlation between the two global measures (CKA linear and RSA Spearman) is somewhat lower, it still shows moderate correlation. Even with this reduced correlation strength, similarity consistency measured using RSA Spearman exhibits the same patterns as CKA linear, particularly regarding the model category we found to be most influential – the training objective (see Fig. 13). This additional analysis suggests that our observations on similarity consistency do not depend on the metric used.
Figure 12: Similarity consistency values measured with different metrics. Left: CKA RBF 0.4 versus CKA linear. Right: RSA Spearman versus CKA linear. Correlations are measured using Pearson correlation. Each point represents the consistency value for a pair of datasets and model sets.
Figure 13: Distribution of similarity consistency values for each model set pair within the training objective category, based on RSA Spearman representational similarities
Appendix F Validating training objective effects through controlled anchor model analysis
To validate our findings regarding the influence of different model categories (Sec. 4.5), we conducted a controlled analysis using six carefully selected anchor models: OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, and ViT-L. These models were chosen to create systematic variations across training objectives (image-text, self-supervised, and supervised learning), while controlling for architecture (convolutional ResNet-50 vs transformer ViT-L) and model size. This controlled setup helps verify whether the patterns observed in our main analysis persist when comparing individual representative models to broader model sets. For each anchor model, we define a single-element model set (e.g., $\Phi_{\text{OpenCLIP RN50}}$) and compute its similarity consistency with different model groups (e.g., ${\bm{R}}(\Phi_{\text{OpenCLIP RN50,}},{\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}})$.
Fig. 14 shows the correlation coefficient distributions across dataset pairs, focusing on training objectives as they emerged as the most influential factor for similarity consistency. The results support our main findings while providing additional insights. Self-supervised models (SimCLR RN50, DINOv2 ViT-L) consistently show the highest median correlations and lowest variations when compared with any anchor model, regardless of the anchor’s training objective. Surprisingly, even image-text (OpenCLIP RN50, OpenCLIP ViT-L) and supervised (ResNet-50, ViT-L) anchor models correlate more strongly with ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$than with their own categories (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$and ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$, respectively). We also observe a notable pattern of weak correlations between image-text anchors and supervised models (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) and vice versa. These patterns remain remarkably consistent across different architectures and model sizes, suggesting that training objectives, rather than architectural choices or model sizes, primarily drive similarity relationships across datasets.
Figure 14: Distribution of Pearson correlation coefficients between anchor models (OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, ViT-L) and models trained with different objectives. Each box shows correlations across all dataset pairs between one anchor model and models from specific sets (e.g., ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$). Results for CKA RBF ($\sigma=0.4$) and CKA linear are separated by dotted lines. Subplot titles are colored according to the anchor model’s subcategory.
Appendix G Distribution correlation coefficients downstream task performance vs. model similarity
In Sec. 4.7, we examined the correlations between representational similarity and classification performance gaps across three datasets from each of the four dataset categories: Natural (multi-domain), Natural (single-domain), Specialized, and Structured. We observed stronger negative correlations for the Natural (single-domain) datasets. Fig. 15 extends this analysis by illustrating the correlation coefficients for all evaluated datasets. We can confirm that natural (single-domain) datasets exhibit the strongest negative correlations overall on this larger set of datasets, with the SVHN dataset being the sole outlier.
Interestingly, ImageNet-1k exhibits the weakest negative correlation between performance gap and representational similarity. We hypothesize that this is due in part to the high diversity of classes in ImageNet-1k, along with its rich set of confounders and contextual cues, which enable networks to approach the classification task in multiple ways. This may be exacerbated by the fact that some of the evaluated networks were pretrained on ImageNet-1k or ImageNet-21k, which could have encouraged the learning of dataset-specific features that perform well on ImageNet but fail to generalize to other datasets.
While our categorization of natural datasets into single-domain and multi-domain revealed an interesting difference in the relationship between downstream task performance and model similarity—especially between the ImageNet-1k and single-domain datasets like Flowers or Pets—the correlation coefficients for these categories still overlap. We hypothesize that our binary sub-categorization of Natural datasets is insufficient for capturing a clear distinction for two reasons: (1) Datasets lie on a spectrum between the two (hypothetical) extremes of encompassing all possible domains and encompassing exactly one. For instance, ImageNet-1k has a relatively rich set of domains, containing 1.000 diverse classes, whereas CIFAR-10 only contains 10 classes of vehicles and animals. (2) Datasets differ in the amount of contextual information (e.g., the object’s background) they offer. We assume that contextual information plays an important role in the number of viable features a dataset classification can be solved with. A large number of viable features means that multiple well-performing but dissimilar strategies may exist. Consequently, multi-domain datasets with limited contextual information may behave more like single-domain datasets. For example, the PASCAL VOC 2007 dataset, when used for single-label classification, makes use of tight bounding-box crops around the object to be classified, reducing the available context. Similarly, in low-resolution datasets like CIFAR-10 and CIFAR-100, contextual information may be reduced due to the loss of background information. While we have not quantified it, we found the Caltech-101 dataset to contain much cleaner object backgrounds than ImageNet-1k, which also suggests reduced contextual information.
Figure 15: Correlation coefficients between downstream task performance and model representational similarity (CKA linear) grouped by dataset categories measured with Pearson correlation.